The prices of inputs K and L are given as $12 per unit and $3 per unit respectively, and a firm operates with the production function

Q=25 K ^0.5 L^0.5

Q=25K0.5L0.5

.

(i) What is the minimum cost of producing 1,250 units of output?


1
Expert's answer
2022-02-13T12:11:23-0500

"Q= 25K^{0.5}L^{0.5}"

"MPK= 25\times 0.5K^{-0.5}L^{0.5}"

"= 12.5K^{-0.5}L^{0.5}"


"MPL= 25\times 0.5K^{0.5}L^{-0.5}"


"= 12.5K^{0.5}L^{-0.5}"


"\frac{MPL}{MPK}=\frac{w}{r}"


"\frac{12.5K^{0.5}L^{-0.5} }{12.5K^{-0.5}L^{0.5}}=\frac{3}{12}"


"\frac{K}{L}=\frac{3}{12}"


L= 4K

K= "\frac{1}{4}L" = 0.25L

Plug values the above in the production function

"1250= 25(0.25L^{0.5})L^{0.25}"

1250= 31.25L

L= 40

1250= 25K"^{0.5}4K^{0.5}"

1250= 100K

K= 12.5


TC= wL+rK

= "(40\times3)+(12\times12.5)= 270"